# Explainable Concept Generation through Vision-Language Preference Learning

### v2

@@ D.3.1 Cumulative rewards
D.3.2 Action Selection Optimization During RLPO Training
@@ D.3.4 C-deletion
D.4 RLPO in sentiment analysis
@@
A list of positive and negative prompts was created, analogous to class images in traditional image classification tasks. Random prompts, similar to those in Fig. 12, were used to simulate random classes. Every word in the prompt, excluding stop words, along with its synonyms, was treated as a concept for this experiment. Synonyms were generated using the Mistral-7B Instruct model, serving a role comparable to the image generation model in image-based settings. We observed that multiple words were identified, along with their influence on the overall prompt, for both classes as shown in Fig. 10 and Fig. 26.
@@
D.5 Effects of Fine-tuning
@@
To demonstrate this experiment, a ResNet50 model, defined in Appendix C.1, was used. We fine-tuned only the final layer of the neural network by setting the learning rate to 0.001 and momentum to 0.9 using the Stochastic Gradient Descent (SGD) optimizer. Maintaining a low learning rate was crucial to preserving high accuracy. RLPO was then applied to the fine-tuned ResNet50, and a shift in the learned concepts was observed, as shown in Fig. 9. This process highlights the model’s sensitivity to fine-tuning and how training on a subset can shift its conceptual interpretation of images.
D.6 Human survey

### v3

@@ C.3.1 Cumulative Rewards
C.3.2 Action Selection Optimization During RLPO Training
@@ C.3.4 C-deletion
C.4 Abstract Concepts
While generating concepts with RLPO, we can observe the progression of output concepts produced by the SD model. As shown in Fig. 25, applying RLPO to the seed prompt “zoo” on the tiger class reveals different levels of abstraction. These abstractions provide insight into what the model prioritizes when identifying a tiger—progressing from a four-legged orange-furred animal to one with black and white stripes, and finally to an orange-furred animal with stripes and whiskers. We obtain concepts at varying levels of abstraction by adjusting $\eta$, though our method currently cannot determine $\eta$ to achieve a specific abstraction level.
Figure 25: Different levels of abstraction for the “Tiger” class on the GoogleNet classifier are illustrated. The generated image starts as a random “zoo” image and gradually transitions to images with tiger-like features. Observe that the seed prompt “zoo” becomes more animal-like at t=10, gains more stripes at t=20, develops tiger-specific colors at t=30, and progressively refines into a tiger image. The model’s prediction also evolves, starting from a random classification of “oxcart” to confidently identifying the generated concept as “tiger”.
C.5 Qualitative Comparison Between With and Without RL Preference Optimization
Figure 26: “Stripes” concept generation from SD models optimized with and without RL
To demonstrate the usefulness of RL based preference optimization, we compared the output generated from the diffusion model fine-tuned using RL with the one fine-tuned iteratively over every seed prompt (brute-force). As shown in Fig. 26, if we don’t use RL to optimize, we see that stripes seed does not converge to a good quality concept compared to the one obtained using RL for the same time budget. The main reason for that is, the RL agent learns over time what trajectories are worth optimizing and drops the less explainable trajectories, highlighting the need for using RL agent while optimization.
C.6 Qualitative Comparison With Other Popular XAI Techniques
We compare the output generated by other popular XAI techniques such as LIME and GradCam with ones generated by RLPO. As shown in Fig. 27, we can see that other methods just explains where the model is looking at whereas our approach also explains what type of features is the model focuses on.
C.7 RLPO in Sentiment Analysis
@@
A list of positive and negative prompts was created, analogous to class images in traditional image classification tasks. Random prompts, similar to those in Fig. 11, were used to simulate random classes. Every word in the prompt, excluding stop words, along with its synonyms, was treated as a concept for this experiment. Synonyms were generated using the Mistral-7B Instruct model, serving a role comparable to the image generation model in image-based settings. We observed that multiple words were identified, along with their influence on the overall prompt, for both classes (positive and negative) as shown in Fig. 9 and Fig. 28.
@@
C.8 Effects of Fine-Tuning
@@
To demonstrate this we conducted an experiment. In this experiment, we choose a pretrained GoogleNet classifier for the Tiger class whose important seed prompts were ‘orange black and white’, ‘orange and black’, and ‘blurry image’ with TCAV scores of 0.66, 0.66, and 0.62, respectively. Out of these seed prompts, ‘orange black and white’ and ‘orange and black’ highlight the tiger pixels while ‘blurry image’ seed prompt highlights the background pixels (see Fig. 12). This tells us that in order to classify a tiger, GoogleNet looks at both the foreground and background. Now, the engineers want the classifier to classify the tiger based on tiger pixels, not its background (note: from the classical Wolfe-Husky example in LIME (Ribeiro et al., 2016), we know the spurious correlation of background).
To this end, we generated 100 tiger images based on concepts related to ‘orange black and white’ and ‘orange and black’ using a separate generative model and fine-tuned our Googlenet model. Running RLPO on this fine-tuned model revealed that the model learned some new concepts such as ‘whiskers’ and also revealed that previous concepts such as ‘orange black and white’ and ‘orange and black’ are now more important with TCAV scores of 1.0 and 1.0, respectively. This means that the classifier is now only looking at tiger pixels, not the background. Dataset samples are shown in Fig. 29.
Figure 29: Dataset samples used to fine-tune GoogleNet classifier on Tiger class.
C.9 Human Survey: Understanding Human Capabilities
@@
C.10 Human Study: Usability
We conducted a human study to evaluate the usefulness of the provided explanations, involving 19 Machine Learning (ML) engineers. Fig. 31 shows the survey used while conducting the human study. This study was approved by the Institutional Review Board of Arizona State University (IRB ID STUDY00021561).
Figure 31: Screenshots from our human survey with sample questions to validate usability.